# Monocular 3D lane detection for Autonomous Driving: Recent Achievements, Challenges, and Outlooks

### v1

@@ II MONOCULDAR 3D LANE DETECTION IN AD
In [107], researchers introducesd the largest real-world 3D lane detection dataset, ONCE-3DLanes dataset, and provides a more comprehensive evaluation metric to rekindle interest in this task in real-world scenarios. Additionally, the paper proposes a method called SALAD, which enables direct generation of 3D lane layouts from front view images without the need for converting the feature map into the bird’s-eye view (BEV), the network architecture of SALAD is shown in Fig. 5. This paper [45] proposes a novel loss function that utilizes the geometry structure prior of lanes in 3D space, enabling stable reconstruction from local to global and providing explicit supervision. It introduces a 2D lane feature extraction module that leverages direct supervision from the top view, ensuring maximum retention of lane structural information, especially in distant areas, the overrall pipeline is shown is Fig. 7. Additionally, the paper presents a task-specific data augmentation method for 3D lane detection to address the imbalanced data distribution of ground slope and camera pose, enhancing generalization on rare cases. Bai et al. proposed CurveFormer [7], a Transformer-based algorithm for 3D lane detection. In this paper, researchers formulate the queries in the decoder layers as a dynamic anchor point set, and utilize a curve cross-attention module to compute the similarity between the queries and image features. Besides, they also introduce a context sampling unit to predict offsets by combining reference features and queries, guiding the learning of sampling offsets. Ai et al. proposed WS-3D-Lane [3], which for the first time proposed a weakly supervised 3D lane detection method using only 2D lane labels, and outperformed the previous 3D-LaneNet [24] method in evaluation. In addition, the author proposes a camera pitch self-calibration method, which can calculate the camera pitch angle online in real time, thereby reducing the error caused by the pitch angle change between the camera and the ground plane caused by the uneven road surface. In BEV-LaneDet [89], the author proposed virtual camera, a novel data prediction processing module, to unify the extrinsic parameters of cameras, and the consistency of data distribution, the authors proposed keypoint representation, a simple yet effective representation of 3D lane structures. In addition, the MLP Based Spatial Transformation Pyramid was proposed, a lightweight structure that realizes the conversion from multi-angle visual features to BEV features. Huang et al. proposed the Anchor3DLane [31] framework, which directly defines anchor points in 3D space and directly regresses 3D lanes from the front view without introducing a top view, as shown in Fig. 6. The authors also proposed a multi-frame extension of Anchor3DLane to exploit well-aligned temporal information and further improve performance. Furthermore, a global optimization method is developed to fine-tune the lane lines by utilizing the lane iso-width property.
@@
Li et al. proposed a method [45] to directly extract top view lane information from front-view images, reducing the structural loss in the 2D lane representation. The overall pipeline of this method is shown in the Fig. 7. In this paper, the authors considered 3D lane detection as a reconstruction problem from 2D images to 3D space. They proposed that explicit imposition of the geometric prior of 3D lanes during the training process is essential for fully utilizing the structural constraints of inter-lane and intra-lane relationships, and for extracting the height information of 3D lanes from the 2D lane representation. The authors analyzed the geometric relationship between 3D lanes and their 2D representation, and proposed an auxiliary loss function based on the geometric structure prior. They also demonstrated that explicit geometric supervision can enhance noise elimination, outlier rejection, and structure preservation for 3D lanes. Bai et al. proposed CurveFormer [7] and CurveFormer++ [8], which are single-stage methods based on Transformer that can directly calculate the parameters of 3D lanes and can bypass the challenging view transformation step. Specifically, they formalized 3D lane detection as a curve propagation problem using curve queries. A 3D lane query is represented by a dynamic and ordered set of anchor points. By using queries with curve representation in the Transformer decoder, the 3D lane detection results are iteratively refined. Additionally, they introduced a curve cross-attention module to calculate the similarities between curve queries and image features. Furthermore, a context sampling module is provided to capture more relevant image features of a curve query, further enhancing the performance of 3D lane detection. Similar to [66], Li et al. proposed GroupLane [48], a row-wise classification-based method for 3D lane detection. GroupLane’s design consists of two groups of convolutional heads, each corresponding to a lane prediction. This grouping decouples the interaction of information among different lanes, reducing optimization difficulty. During training, predictions are matched with lane labels using a single-win one-to-one matching (SOM) strategy, which assigns predictions to the most suitable labels for loss computation. To address the challenge posed by the inevitable depth ambiguity in monocular images causing misalignment between the constructed surrogate feature map and the original image for lane detection, Luo et al. proposed a novel LATR model [53]. This is an end-to-end 3D lane detector that uses 3D-aware front-view features without the need for transformed view representation. Specifically, LATR detects 3D lanes through cross-attention based on query and key-value pairs constructed using a lane-aware query generator and dynamic 3D ground positional embedding. On one hand, each query is generated based on 2D lane-aware features and adopts a hybrid embedding to enhance lane information. On the other hand, 3D spatial information is injected as positional embedding from an iteratively updated 3D ground plane. To address the issue of inaccurate view transformation caused by the neglect of road height changes in the process of transforming image-view features into bird’s-eye view using inverse perspective mapping (IPM), Chen et al. proposed an efficient transformer for 3D lane detection [15]. Unlike conventional transformers, the model includes a decomposed cross-attention mechanism that can simultaneously learn lane and bird’s-eye view representations. This approach allows for more accurate view transformation compared to IPM-based methods and is more efficient. Previous studies assumed that all lanes are on a flat ground plane. However, Kim et al. argued that algorithms based on this assumption have difficulty in detecting various lanes in actual driving environments and proposed a new algorithm, D-3DLD [40]. Unlike previous methods, this approach expands rich contextual features from the image domain to 3D space by utilizing depth-aware voxel mapping. Additionally, the method determines 3D lanes based on voxelized features. The authors designed a new lane representation combined with uncertainties and predicted the confidence intervals of 3D lane points using Laplace loss. Li and colleagues proposed a lightweight method [46] that uses MobileNet [29] as the backbone network to reduce the demand for computational resources. The proposed method includes the following three stages. First, the MobileNet model is used to generate multi-scale front-view features from a single RGB image. Then, a perspective transformer calculates the bird’s eye view (BEV) features from the front-view features. Finally, two convolutional neural networks are used to predict the 2D and 3D coordinates and their respective lane types. In paper [26], Han et al. argued that the curve-based lane representation may not be well-suited for many irregular lane lines in real-world scenarios, which can lead to performance gaps compared to indirect representations such as segmentation-based or point-based methods. In this paper, the authors proposed a new lane detection method, which can be decomposed into two parts: curve modeling and ground height regression. Specifically, a parameterized curve is used to represent lanes in the bird’s eye view space to reflect the original distribution of lanes. For the second part, since ground heights are determined by natural factors such as road conditions and are not comprehensive, the ground heights are regressed separately from the curve modeling. Additionally, the authors have designed a new framework and a series of losses to unify the 2D and 3D lane detection tasks, guiding the optimization of models with or without 3D lane labels.
@@
An intuitive summary of these methods is shown in Table II, which includes a description of the methods, the datasets used, the open-source status, and the network architecture.
@@ III PERFORMANCE EVALUATION OF 3D LANE DETECTION
This section will discuss the performance evaluation of monocular 3D lane detection models. Herein, we explain the evaluation metrics, different types of objective functions, analyze the computational complexity, and finally provide quantitative comparisons of various models. The nomenclature of the used variables is given in Table LABEL:tab:variables.
@@ III-B1 MSE loss
MSE loss is one of the most commonly used loss functions, which calculates the squared difference between the model’s predicted values and the true values, and then takes the average. Its mathematical expression is:
@@ III-B2 MAE loss
MAE loss is another commonly used loss function, which calculates the absolute difference between the model’s predicted values and the true values, and then takes the average. Its mathematical expression is:
@@ III-B5 Binary Cross-Entropy Loss
Binary Cross-Entropy loss is commonly used for training binary classification tasks, aiming to minimize the loss function to improve the accuracy of the model’s predictions on binary classification samples. It is widely applied in deep learning for tasks such as image classification, text classification, and more.. Its mathematical expression is:
@@ III-B7 IoU Loss
Different methods utilize specific loss functions differently, but essentially, most are variations or combinations of the basic loss functions mentioned above. Additionally, the Hungarian algorithm [100] is often used to match predicted lanes with ground truth lanes.
@@ III-C Quantitative Analysis of Monocular 3D Lane Detection Models
Additionally, computational efficiency is taken into account by reporting the Frames Per Second (FPS) achievable by each method during the inference process. The overall runtime of these models is reported in Table IV,V and VI. In some cases, the time complexity indicators of the models are reported from their methods, while in other cases, we ran the reported models from their publicly available repositories using our experimental resources. The system’s CPU configuration includes an Intel(R) Core i9-12900K CPU processor running on the Ubuntu 20.04 operating system, while the GPU used in the experiments is an NVIDIA GeForce RTX 3080Ti GPU with 12 GB of graphics memory, In the table, we specified the hardware used for inference for each method.
@@ IV DATASETS
In vision tasks based on deep learning, another equally important component is the dataset. In this section, we will introduce the datasets currently available for monocular 3D lane line detection tasks. Some of these datasets are open-sourced and popular, frequently used by the community, while others are only described in papers and not open-sourced. We have compiled an intuitive table of all existing monocular 3D lane line detection datasets, as shown in Table III.
@@ IV-A Apollo 3D Lane Synthetic Dataset
The Apollo 3D Lane Synthetic Dataset [25] is a robust synthetic dataset comprising 10,500 frames of high-resolution 1080 × 1920 monocular RGB images, built using the Unity 3D engine. Each frame is accompanied by corresponding 3D lane labels and camera pitch data. The dataset, based on Silicon Valley in the United States, spans a wide array of environments, including highways, urban areas, residential neighborhoods, and downtown settings. The dataset’s images encapsulate a broad range of daytime and weather conditions, various traffic/obstacle scenarios, and different road surface qualities, thereby providing the dataset with a high degree of diversity and realism.
@@
Network Architecture   Need BEV Trans.?
[24]
@@
2018 Combining novel network architecture, IPM, and anchor-based lane representation to achieve accurate 3D lane detection. 3D-Lanes, Synthetic-3D-Lanes
   N/A
   CNN                 ✔
[25]
@@
2020 Using a two-stage framework that decouples the learning of the image segmentation subnetwork and the geometry encoding subnetwork, reducing the amount of 3D lane labels required for training.
Apollo Synthetic.
   link
   CNN                 ✔
[17]
@@
2020 Introducing uncertainty estimation, utilizing a semi-local, Bird’s Eye View (BEV), tile representation to break down lanes into simple segments, and subsequently clustering them into complete lanes.
Synthetic-3D-Lanes, 3D-lanes.
   N/A
   CNN                 ✔
[16]
@@
2020 An anchor-free method, which can detect complex lane topologies such as splits, merges, short lanes, and perpendicular lanes.
Synthetic-3D-Lanes, Real-3D-lanes
   N/A
   CNN                 ✔
[52]
@@
2021 Presenting a two-stage framework, leveraging auxiliary 3D lane tasks and geometry constraints in the first stage, followed by accurate 3D lane prediction in the second stage using estimated pose.
Apollo Synthetic.
   link
   CNN                 ✔
[12]
@@
2022 The paper introduces a novel Transformer-based spatial feature transformation module and presents OpenLane, a large-scale real-world 3D lane dataset.
Apollo Synthetic, OpenLane.
   link
Transformer           ✔
[107]
@@
2022 The paper introduces the ONCE-3DLanes dataset and proposes SALAD, an extrinsic-free, anchor-free method for regressing 3D lane coordinates in image view.
ONCE-3DLanes.
   link
   CNN                 ✘
[33]
@@
2022 The paper proposes Att-Gen-LaneNet, a two-stage network for robust 3D lane detection. It integrates the Efficient Channel Attention (ECA) and Convolutional Block Attention Module (CBAM) to improve segmentation and generalization.
Apollo Synthetic.
   N/A
   CNN                 ✔
[45]
@@
2022 The paper proposes an approach for monocular 3D lane detection by leveraging the geometry structure and explicit supervision based on the structure prior.
Apollo Synthetic.
   N/A
   CNN                 ✔
[46]
@@
2022 The method comprises three stages: generating multiscale front-view features, calculating bird’s eye view (BEV) features, and predicting 2D and 3D lanes.
OpenLane.
   N/A
   CNN                 ✔
[7]
@@
2022 The method formulates 3D lane detection as a curve propagation problem, enhancing performance by leveraging curve queries, a curve cross-attention module, and a context sampling module to capture relevant image features.
Apollo Synthetic, OpenLane.
   N/A
Transformer           ✘
[3]
@@
2022 The paper introduces a method training 3D lanes with only 2D lane labels, leveraging assumptions of constant lane width and equal height on adjacent lanes for indirect supervision of 3D lane heights.
Apollo Synthetic, ONCE-3DLanes.
   N/A
   CNN                 ✔
[90]
@@
2022 The paper introduces a Virtual Camera for spatial relationship consistency, a Key-Points Representation module for diverse 3D lane structures, and a Spatial Transformation Pyramid for front-view to bird’s eye view feature transformation.
Apollo Synthetic, OpenLane.
   link
   CNN                 ✔
[95]
@@
2022 Utilizing multi-frame input and removing the dependency on camera pose, the model integrates a multi-frame pre-alignment layer, spatio-temporal attention module, and 3DLane IOULoss.
OpenLane, ONCE-3DLanes.
   N/A
   CNN                 ✔
[31]
@@
2023 The paper introduces Anchor3DLane, a BEV-free method predicting 3D lanes directly from front-viewed (FV) representations using defined 3D lane anchors.
Apollo Synthetic, OpenLane, ONCE-3DLanes.
   link
   CNN                 ✘
[40]
@@
2023 The paper introduces a method that employs depth-aware voxel mapping, extends contextual features to 3D space, and predicts confidence intervals using Laplace loss.
LLAMAS-3DLD, Pandaset-3DLD, Apollo Synthetic.
   N/A
   CNN                 ✔
[15]
@@
2023 The paper proposes an efficient transformer for 3D lane detection using a decomposed cross-attention mechanism to learn lane and bird’s-eye-view (BEV) representations concurrently.
OpenLane, ONCE-3DLanes.
   N/A
Transformer           ✘
[48]
@@
2023 Similar to Paper [66], this is the first work on 3D lane detection utilizing fully convolutional heads based on row-wise classification.
OpenLane, ONCE-3DLanes, OpenLane-Huawei.
   N/A
   CNN                 ✔
[53]
@@
2023 The paper introduces LATR, an end-to-end 3D lane detection model utilizing 3D-aware front-view features and cross-attention mechanism with lane-aware query generator.
OpenLane, Apollo Synthetic, ONCE-3DLanes.
   link
Transformer           ✘
[26]
@@
2023 The paper introduces a novel lane detection method, dividing it into curve modeling and ground height regression, within a unified framework for both 2D and 3D lane detection.
OpenLane, ONCE-3DLanes, TuSimple, CULane.
   N/A
   CNN                 ✘
[8]
@@
2024 The approach models the 3D detection task as a curve propagation problem, where each lane is represented by a curve query with a dynamic and ordered anchor point set.
OpenLane, ONCE-3DLanes.
   N/A
Transformer           ✘
@@
The dataset is segmented into three distinct scene categories: Balanced Scenes, Rarely Observed Scenes, and Scenes with Visual Variations. Balanced Scenes serve as a comprehensive and unbiased dataset for benchmarking standard driving scenarios. Rarely Observed Scenes are used for testing an algorithm’s adaptability to rarely encountered situations involving complex urban maps with drastic elevation changes and sharp turns. Scenes with Visual Variations are designed to assess how algorithms perform under different illumination conditions by excluding specific daytime periods during training and focusing on them during testing. The camera’s intrinsic parameters in the dataset are fixed, with the camera heights ranging between 1.4 to 1.8 meters and pitch angles spanning from 0 to 10 degrees.

### v2

@@ II MONOCULDAR 3D LANE DETECTION IN AD
In [yan2022once], researchers introducesd the largest real-world 3D lane detection dataset, ONCE-3DLanes dataset, and provides a more comprehensive evaluation metric to rekindle interest in this task in real-world scenarios. Additionally, this paper proposes a method called SALAD, which enables direct generation of 3D lane layouts from front view images without the need for converting the feature map into the bird’s-eye view (BEV), the network architecture of SALAD is shown in Fig. 5. This paper [li2022reconstruct] proposes a novel loss function that utilizes the geometry structure prior of lanes in 3D space, enabling stable reconstruction from local to global and providing explicit supervision. It introduces a 2D lane feature extraction module that leverages direct supervision from the top view, ensuring maximum retention of lane structural information, especially in distant areas, the overrall pipeline is shown is Fig. 7. Additionally, the paper presents a task-specific data augmentation method for 3D lane detection to address the imbalanced data distribution of ground slope and camera pose, enhancing generalization on rare cases. Bai et al. proposed CurveFormer [bai2022curveformer], a Transformer-based algorithm for 3D lane detection. In this paper, researchers formulate the queries in the decoder layers as a dynamic anchor point set, and utilize a curve cross-attention module to compute the similarity between the queries and image features. Besides, they also introduce a context sampling unit to predict offsets by combining reference features and queries, guiding the learning of sampling offsets. Ai et al. proposed WS-3D-Lane [ai2022ws], which for the first time proposed a weakly supervised 3D lane detection method using only 2D lane labels, and outperformed the previous 3D-LaneNet [garnett20193d] method in evaluation. In addition, the author proposed a camera pitch self-calibration method, which can calculate the camera pitch angle online in real time, thereby reducing the error caused by the pitch angle change between the camera and the ground plane caused by the uneven road surface. In BEV-LaneDet [wang2022bev], the author introduced the concept of a virtual camera, a novel module for data prediction processing, aiming to standardize the extrinsic parameters of cameras and ensure the consistency of data distribution. Furthermore, they proposed the use of keypoint representation as a simple yet effective method for representing 3D lane structures. Additionally, they presented the MLP Based Spatial Transformation Pyramid, a lightweight architecture designed to convert multi-angle visual features into BEV features. Huang et al. proposed the Anchor3DLane [huang2023anchor3dlane] framework, which directly defines anchor points in 3D space and directly regresses 3D lanes from the front view without introducing a top view, as shown in Fig. 6. The authors also proposed a multi-frame extension of Anchor3DLane to exploit well-aligned temporal information and further improve performance. Furthermore, a global optimization method is developed to fine-tune the lane lines by utilizing the lane iso-width property.
@@
Li et al. proposed a method [li2022reconstruct] to directly extract top view lane information from front-view images, reducing the structural loss in the 2D lane representation. The overall pipeline of this method is shown in the Fig. 7. In this paper, the authors considered 3D lane detection as a reconstruction problem from 2D images to 3D space. They proposed that explicit imposition of the geometric prior of 3D lanes during the training process is essential for fully utilizing the structural constraints of inter-lane and intra-lane relationships, and for extracting the height information of 3D lanes from the 2D lane representation. The authors analyzed the geometric relationship between 3D lanes and their 2D representation, and proposed an auxiliary loss function based on the geometric structure prior. They also demonstrated that explicit geometric supervision can enhance noise elimination, outlier rejection, and structure preservation for 3D lanes. Bai et al. proposed CurveFormer [bai2022curveformer] and CurveFormer++ [bai2024curveformer++], which are single-stage methods based on Transformer that can directly calculate the parameters of 3D lanes and can bypass the challenging view transformation step. Specifically, they formalized 3D lane detection as a curve propagation problem using curve queries. A 3D lane query is represented by a dynamic and ordered set of anchor points. By using queries with curve representation in the Transformer decoder, the 3D lane detection results are iteratively refined. Additionally, they introduced a curve cross-attention module to calculate the similarities between curve queries and image features. Furthermore, a context sampling module is provided to capture more relevant image features of a curve query, further enhancing the performance of 3D lane detection. Similar to [qin2020ultra], Li et al. proposed GroupLane [li2023grouplane], a row-wise classification-based method for 3D lane detection. GroupLane’s design consists of two groups of convolutional heads, each corresponding to a lane prediction. This grouping decouples the interaction of information among different lanes, reducing optimization difficulty. During training, predictions are matched with lane labels using a single-win one-to-one matching (SOM) strategy, which assigns predictions to the most suitable labels for loss computation. To address the challenge posed by the inevitable depth ambiguity in monocular images causing misalignment between the constructed surrogate feature map and the original image for lane detection, Luo et al. proposed a novel LATR model [luo2023latr]. This is an end-to-end 3D lane detector that uses 3D-aware front-view features without the need for transformed view representation. Specifically, LATR detects 3D lanes through cross-attention based on query and key-value pairs constructed using a lane-aware query generator and dynamic 3D ground positional embedding. On one hand, each query is generated based on 2D lane-aware features and adopts a hybrid embedding to enhance lane information. On the other hand, 3D spatial information is injected as positional embedding from an iteratively updated 3D ground plane. To address the issue of inaccurate view transformation caused by the neglect of road height changes in the process of transforming image-view features into bird’s-eye view using inverse perspective mapping (IPM), Chen et al. proposed an efficient transformer for 3D lane detection [chen2023efficient]. Unlike conventional transformers, the model includes a decomposed cross-attention mechanism that can simultaneously learn lane and bird’s-eye view representations. This approach allows for more accurate view transformation compared to IPM-based methods and is more efficient. Previous studies assumed that all lanes are on a flat ground plane. However, Kim et al. argued that algorithms based on this assumption have difficulty in detecting various lanes in actual driving environments and proposed a new algorithm, D-3DLD [kim2023d]. Unlike previous methods, this approach expands rich contextual features from the image domain to 3D space by utilizing depth-aware voxel mapping. Additionally, the method determines 3D lanes based on voxelized features. The authors designed a new lane representation combined with uncertainties and predicted the confidence intervals of 3D lane points using Laplace loss. Li et al. proposed a lightweight method [li2022perspective] that uses MobileNet [howard2017mobilenets] as the backbone network to reduce the demand for computational resources. The proposed method includes the following three stages. First, the MobileNet model is used to generate multi-scale front-view features from a single RGB image. Then, a perspective transformer calculates the bird’s eye view (BEV) features from the front-view features. Finally, two convolutional neural networks are used to predict the 2D and 3D coordinates and their respective lane types. In paper [han2023decoupling], Han et al. argued that the curve-based lane representation may not be well-suited for many irregular lane lines in real-world scenarios, which can lead to performance gaps compared to indirect representations such as segmentation-based or point-based methods. In this paper, the authors proposed a new lane detection method, which can be decomposed into two parts: curve modeling and ground height regression. Specifically, a parameterized curve is used to represent lanes in the bird’s eye view space to reflect the original distribution of lanes. For the second part, since ground heights are determined by natural factors such as road conditions and are not comprehensive, the ground heights are regressed separately from the curve modeling. Additionally, the authors have designed a new framework and a series of losses to unify the 2D and 3D lane detection tasks, guiding the optimization of models with or without 3D lane labels.
@@
An intuitive summary of these methods is shown in Table LABEL:tab:overviw, which includes a description of the methods, the datasets used, the open-source status, and the network architecture.
@@ III PERFORMANCE EVALUATION OF 3D LANE DETECTION
This section will discuss the performance evaluation of monocular 3D lane detection models. Herein, we explain the evaluation metrics, different types of objective functions, analyze the computational complexity, and finally provide quantitative comparisons of various models. The nomenclature of the used variables is given in Table I.
@@ III-B1 MSE loss
This loss is one of the most commonly used loss functions, which calculates the squared difference between the model’s predicted values and the true values, and then takes the average. Its mathematical expression is:
@@ III-B2 MAE loss
This loss is another commonly used loss function, which calculates the absolute difference between the model’s predicted values and the true values, and then takes the average. Its mathematical expression is:
@@ III-B5 Binary Cross-Entropy Loss
Binary Cross-Entropy loss is commonly used for training binary classification tasks, aiming to minimize the loss function to improve the accuracy of the model’s predictions on binary classification samples. It is widely applied in deep learning for tasks such as image classification, text classification, and segmentation. Its mathematical expression is:
@@ III-B7 IoU Loss
Different methods utilize specific loss functions differently, but essentially, most are variations or combinations of the basic loss functions mentioned above. Additionally, the Hungarian algorithm [enwiki:1212340485] is often used to match predicted lanes with ground truth lanes.
@@ III-C Quantitative Analysis of Monocular 3D Lane Detection Models
Additionally, computational efficiency is taken into account by reporting the Frames Per Second (FPS) achievable by each method during the inference process. The overall runtime of these models is reported in Table IV,V and VI. In some papers, the inference time of the algorithms and their corresponding hardware platforms are reported, which we directly utilize. However, in some other papers, the inference time of the algorithms is not shown, so we conduct our own experiments to test the inference time on our experimental platform. Our experimental platform’s CPU configuration includes an Intel(R) Core i9-12900K CPU processor running on the Ubuntu 20.04 operating system, while the GPU used in the experiments is an NVIDIA GeForce RTX 3080Ti GPU with 12 GB of graphics memory, In the above tables, we specified the hardware used for inference for each method.
@@ IV DATASETS
In vision tasks based on deep learning, another equally important component is the dataset. In this section, we will introduce the datasets currently available for monocular 3D lane line detection tasks. Some of these datasets are open-sourced and popular, frequently used by the community, while others are only described in papers and not open-sourced. Whether open-source or proprietary datasets, in order to provide a more intuitive overview of these datasets, we have compiled a detailed table showcasing all existing monocular 3D lane line detection datasets, as presented in Table III.
@@ IV-A Apollo 3D Lane Synthetic Dataset
The Apollo 3D Lane Synthetic Dataset [guo2020gen] is a robust synthetic dataset comprising 10,500 frames of high-resolution 1080 × 1920 monocular RGB images, built using the Unity 3D engine. Each frame is accompanied by corresponding 3D lane labels and camera pitch data. It based on Silicon Valley in the United States, spans a wide array of environments, including highways, urban areas, residential neighborhoods, and downtown settings.
@@
Ref. Year Methodology Dataset Code Network Architecture   Need BEV Trans.?
@@
[garnett20193d] 2018 Combining novel network architecture, IPM, and anchor-based lane representation to achieve accurate 3D lane detection. 3D-Lanes, Synthetic-3D-Lanes N/A CNN                 ✔
@@
[guo2020gen] 2020 Using a two-stage framework that decouples the learning of the image segmentation subnetwork and the geometry encoding subnetwork, reducing the amount of 3D lane labels required for training. Apollo Synthetic. link CNN                 ✔
@@
[efrat2020semi] 2020 Introducing uncertainty estimation, utilizing a semi-local, Bird’s Eye View (BEV), tile representation to break down lanes into simple segments, and subsequently clustering them into complete lanes. Synthetic-3D-Lanes, 3D-lanes. N/A CNN                 ✔
@@
[efrat20203d] 2020 An anchor-free method, which can detect complex lane topologies such as splits, merges, short lanes, and perpendicular lanes. Synthetic-3D-Lanes, Real-3D-lanes N/A CNN                 ✔
@@
[liu2022learning] 2021 Presenting a two-stage framework, leveraging auxiliary 3D lane tasks and geometry constraints in the first stage, followed by accurate 3D lane prediction in the second stage using estimated pose. Apollo Synthetic. link CNN                 ✔
@@
[chen2022persformer] 2022 This paper introduces a novel Transformer-based spatial feature transformation module and presents OpenLane, a large-scale real-world 3D lane dataset. Apollo Synthetic, OpenLane. link Transformer           ✔
@@
[yan2022once] 2022 This paper introduces the ONCE-3DLanes dataset and proposes SALAD, an extrinsic-free, anchor-free method for regressing 3D lane coordinates in image view. ONCE-3DLanes. link CNN                 ✘
@@
[jiang2022robust] 2022 This paper proposes Att-Gen-LaneNet, a two-stage network for robust 3D lane detection. It integrates the Efficient Channel Attention (ECA) and Convolutional Block Attention Module (CBAM) to improve segmentation and generalization. Apollo Synthetic. N/A CNN                 ✔
@@
[li2022reconstruct] 2022 This paper proposes an approach for monocular 3D lane detection by leveraging the geometry structure and explicit supervision based on the structure prior. Apollo Synthetic. N/A CNN                 ✔
@@
[li2022perspective] 2022 The method comprises three stages: generating multiscale front-view features, calculating bird’s eye view (BEV) features, and predicting 2D and 3D lanes. OpenLane. N/A CNN                 ✔
@@
[bai2022curveformer] 2022 The method formulates 3D lane detection as a curve propagation problem, enhancing performance by leveraging curve queries, a curve cross-attention module, and a context sampling module to capture relevant image features. Apollo Synthetic, OpenLane. N/A Transformer           ✘
@@
[ai2022ws] 2022 This paper introduces a method training 3D lanes with only 2D lane labels, leveraging assumptions of constant lane width and equal height on adjacent lanes for indirect supervision of 3D lane heights. Apollo Synthetic, ONCE-3DLanes. N/A CNN                 ✔
@@
[wang2023bevlanedet] 2022 This paper introduces a Virtual Camera for spatial relationship consistency, a Key-Points Representation module for diverse 3D lane structures, and a Spatial Transformation Pyramid for front-view to bird’s eye view feature transformation. Apollo Synthetic, OpenLane. link CNN                 ✔
@@
[wang2022spatio] 2022 Utilizing multi-frame input and removing the dependency on camera pose, the model integrates a multi-frame pre-alignment layer, spatio-temporal attention module, and 3DLane IOULoss. OpenLane, ONCE-3DLanes. N/A CNN                 ✔
@@
[huang2023anchor3dlane] 2023 This paper introduces Anchor3DLane, a BEV-free method predicting 3D lanes directly from front-viewed (FV) representations using defined 3D lane anchors. Apollo Synthetic, OpenLane, ONCE-3DLanes. link CNN                 ✘
@@
[kim2023d] 2023 This paper introduces a method that employs depth-aware voxel mapping, extends contextual features to 3D space, and predicts confidence intervals using Laplace loss. LLAMAS-3DLD, Pandaset-3DLD, Apollo Synthetic. N/A CNN                 ✔
@@
[chen2023efficient] 2023 This paper proposes an efficient transformer for 3D lane detection using a decomposed cross-attention mechanism to learn lane and bird’s-eye-view (BEV) representations concurrently. OpenLane, ONCE-3DLanes. N/A Transformer           ✘
@@
[li2023grouplane] 2023 Similar to Paper [qin2020ultra], this is the first work on 3D lane detection utilizing fully convolutional heads based on row-wise classification. OpenLane, ONCE-3DLanes, OpenLane-Huawei. N/A CNN                 ✔
@@
[luo2023latr] 2023 This paper introduces LATR, an end-to-end 3D lane detection model utilizing 3D-aware front-view features and cross-attention mechanism with lane-aware query generator. OpenLane, Apollo Synthetic, ONCE-3DLanes. link Transformer           ✘
@@
[han2023decoupling] 2023 This paper introduces a novel lane detection method, dividing it into curve modeling and ground height regression, within a unified framework for both 2D and 3D lane detection. OpenLane, ONCE-3DLanes, TuSimple, CULane. N/A CNN                 ✘
@@
[bai2024curveformer++] 2024 The approach models the 3D detection task as a curve propagation problem, where each lane is represented by a curve query with a dynamic and ordered anchor point set. OpenLane, ONCE-3DLanes. N/A Transformer           ✘
@@
The dataset’s images encapsulate a broad range of daytime and weather conditions, various traffic/obstacle scenarios, and different road surface qualities, thereby providing the dataset with a high degree of diversity and realism. The dataset is segmented into three distinct scene categories: Balanced Scenes, Rarely Observed Scenes, and Scenes with Visual Variations. Balanced Scenes serve as a comprehensive and unbiased dataset for benchmarking standard driving scenarios. Rarely Observed Scenes are used for testing an algorithm’s adaptability to rarely encountered situations involving complex urban maps with drastic elevation changes and sharp turns. Scenes with Visual Variations are designed to assess how algorithms perform under different illumination conditions by excluding specific daytime periods during training and focusing on them during testing. The camera’s intrinsic parameters in the dataset are fixed, with the camera heights ranging between 1.4 to 1.8 meters and pitch angles spanning from 0 to 10 degrees.